OpenStudy (anonymous):

-8 - 8 = 0 :) Right !

OpenStudy (akshay_budhkar):

-8-8=-16 =8+8=0

OpenStudy (savvy):

no man....-8-8 = -16

OpenStudy (savvy):

-8-8 = (-8) + (-8) = -(8+8) = -16

OpenStudy (anonymous):

Negative 8 plus positve 8 = 0

OpenStudy (phi):

-8+8=0

OpenStudy (savvy):

@smoove yes that is right...

OpenStudy (muhammad_nauman_umair):

is-16

OpenStudy (anonymous):

thas wrong..its -16